# Generalized quantum Zernike Hamiltonians: Polynomial Higgs-type algebras and algebraic derivation of the spectrum

### v1

@@ 4 The cubic quantum Zernike Hamiltonian
In this section we study the quantum symmetries and deduce the possible spectra for the cubic Hamiltonian ^$\hat{\mathcal{H}}_{3}$ (2.3) depending on three arbitrary parameters $\gamma_{1}$, $\gamma_{2}$ and $\gamma_{3}$. It is worthy to be observed that the cubic $\gamma_{3}$-term in ^$\hat{\mathcal{H}}_{3}$ can be interpreted as a superintegrable perturbation of the quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ in (2.2). Indeed, we require that all the expressions for ^$\hat{\mathcal{H}}_{2}$ in (2.3) obtained in the previous section are recovered whenever $\gamma_{3}$ is put equal to zero.
The two quantum symmetries (2.6) of ^$\hat{\mathcal{H}}_{3}$, that are of third-order in the momenta, are obtained after some computation, and adopt the form
@@
The number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ are defined exactly as in (3.4), and satisfy the commutation relations (2.9) with
@@
The symmetry algebra of the cubic quantum Hamiltonian ^$\hat{\mathcal{H}}_{3}$ hence corresponds to a fifth-order polynomial Higgs-type algebra $\mathfrak{sl}^{(5)}(2,\mathbb{R})$ (see [3] for the classical picture).
@@ 4.1 Spectrum of the cubic system
(i) The set of equations (2.14) coming from (4.1) yields ten types of solutions for the representation-dependent constant $u=u({n})$ and the spectrum $E=E({n})$ of the Hamiltonian ^$\hat{\mathcal{H}}_{3}$ in (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$. They lead to the structure function $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}=\{1,\dots,{n}\}$ via the factorization (4.1).
(ii) The requirement that the solutions are well-defined for any values of $\gamma_{1}$ and $\gamma_{2}$ for $\gamma_{3}=0$ reduces the ten types of solutions to only two, given by
@@
To illustrate the situation with the seven discarded solutions for the limit $\gamma_{3}\to 0$, let us consider, for instance, one of them given by
@@
The two consistent types of solutions given in proposition 2 constitute a cubic extension of those presented in proposition 1 and displayed in Table 1. We observe that types III and IV do not admit a cubic generalization, implying that they are specific solutions for the quadratic Hamiltonian (2.3) with $N=2$. The spectrum corresponding to types I and II takes real values whenever the coefficients $\gamma_{1}$ and $\gamma_{3}$ are pure imaginary numbers, while $\gamma_{2}$ is real. By considering (1.5), as before, we denote
@@
to be compared with (3.7). Furthermore, if we set $\alpha=-1$ and $\beta=-2$ and identify ${n}$ with the principal quantum number $\mathfrak{n}$, we obtain two possible superintegrable perturbations of the spectrum (3.8) of the Zernike Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ in (1.1), reading as
@@ 4.2 Perturbations of the Euclidean, spherical and hyperbolic oscillators
Following the analysis in section 3.2, we can interpret the results of proposition 2 as superintegrable cubic perturbations of the isotropic oscillator on $\mathbf{E}^{2}$, $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$. To this end, let us consider the solution of type I (4.4) (eliminating the index ‘I’), again set $\beta=-2$ and $\alpha=-\kappa$, but now keeping an arbitrary real parameter $\mu$, which characterizes the perturbation. Thus, we have six possibilities, according to the curvature $\kappa$ and the two signs of $\mu$.
For the flat Euclidean oscillator with $\kappa=0$ we start with a negative value for the perturbation parameter $\mu=-|\mu|$ obtaining that
@@
Hence, the spectrum is cubic, unbounded and the energy between two consecutive states grows quadratically with ${n}$. It is worth noting that the behaviour of the spectrum (4.5) is quite similar to that of the spherical or Higgs oscillator (3.11) and (3.12), so we can say that it is a cubic spherical perturbation of the Euclidean oscillator. In figure 3 we plot the first energies (4.5) for three negative values of $\mu$ together with those of the non-perturbed Euclidean oscillator (3.10) (to be compared with figure 1).
@@
We require a positive spectrum and positive differences of two consecutive energies. As for the hyperbolic oscillator (see (3.15) and (3.16)) such conditions yield a maximum value $n_{\rm max}$ for the positive integer $n\geq 1$ which is the greatest integer number that satisfies the following constraint
$21\leq n_{\rm max}<\frac{1}{2}+\frac{1}{6}\sqrt{\frac{24}{\mu}-3}\quad\Longrightarrow\quad\mu<$ 2 . (4.7)
Here, the perturbation parameter can be expressed in terms of a fixed value of $n_{\rm max}$ in the form
$\begin{array}[]{ll}n_{\rm max}=1\!\!:&\quad\displaystyle\frac{2}{7}\leq\mu<2.\\[10.0pt] n_{\rm max}>1\!\!:&\quad\displaystyle\frac{2}{1+3n_{\rm max}(n_{\rm max}+1)}\leq\mu<\frac{2}{1+3n_{\rm max}(n_{\rm max}-1)}.\end{array}$
In this respect, we can say that (4.6) and (4.7) determine a cubic hyperbolic perturbation of the Euclidean oscillator that leads to a positive finite discrete spectrum with ${n}\in\{1,\dots,n_{\rm max}\}$. We represent the discrete spectrum for three values of $n_{\rm max}$ (so for $\mu$) in figure 4, which are lower than the Euclidean ones (3.10), and worthy of comparison with figure 2.
@@
[…]
Now we consider the sphere $\mathbf{S}^{2}$ with positive curvature $\kappa>0$ and a negative parameter $\mu=-|\mu|$. Thus we arrive at
@@
$E(n)=2n+\kappa n^{2}-\mu n^{3},\qquad E(n+1)-E(n)=2+\kappa(2n+1)-\mu(3n^{2}+3n+1),$ (4.8)
and the requirements for getting a positive spectrum and positive differences of two consecutive energies imply that there is a maximum value $n_{\rm max}$ for the positive integer $n\geq 1$ that has to satisfy the inequality
$1\leq n_{\rm max}<\frac{1}{2}+\frac{\kappa}{3\mu}+\frac{1}{6}\sqrt{\frac{4\kappa^{2}+24\mu}{\mu^{2}}-3}\quad\Longrightarrow\quad\mu<2+\kappa.$ (4.9)
From it, the perturbation parameter $\mu$ can be written in terms of a chosen value for $n_{\rm max}$ as
$\begin{array}[]{ll}n_{\rm max}=1\!\!:&\quad\displaystyle\frac{2+3\kappa}{7}\leq\mu<2+\kappa.\\[10.0pt] n_{\rm max}>1\!\!:&\quad\displaystyle\frac{2+\kappa(2n_{\rm max}+1)}{1+3n_{\rm max}(n_{\rm max}+1)}\leq\mu<\frac{2+\kappa(2n_{\rm max}-1)}{1+3n_{\rm max}(n_{\rm max}-1)}.\end{array}$
@@
Likewise, both types of perturbations can be applied to the oscillator on the hyperbolic space $\mathbf{H}^{2}$ with negative curvature $\kappa=-|\kappa|<0$ (see (3.13)–(3.15)), namely
@@
A similar analysis as above would be necessary for each type of perturbation. However, the results are quite cumbersome, as there are several possibilities for $n_{\rm max}$ depending on $|\kappa|$ and $\mu$. For brevity in the exposition, these explicit computations are omitted.
@@
5 The quartic quantum Zernike Hamiltonian
We proceed similarly to the two previous sections, first giving the quantum symmetries and the possible spectra for the quartic Hamiltonian ^$\hat{\mathcal{H}}_{4}$ in (2.3), depending on four arbitrary parameters $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$ and $\gamma_{4}$, and imposing that the quantum quadratic Hamiltonian ^$\hat{\mathcal{H}}_{2}$ in (2.3) is recovered for the vanishing of $\gamma_{3}$ and $\gamma_{4}$.
The two quantum symmetries (2.6) of ^$\hat{\mathcal{H}}_{4}$, now of fourth-order in the momenta, read
and verify the algebraic dependence relation
^$\hat{\mathcal{H}}_{4}={\hat{I}}_{4}+{\hat{I}}_{4}^{\prime}-4\gamma_{4}\hat{\mathcal{C}}^{2}+\gamma_{4}\hat{\mathcal{C}}^{4}.$
The number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ are now defined by
$\begin{split}\hat{\mathcal{K}}&:=\frac{1}{2}\hat{\mathcal{K}}_{1},\\ \hat{\mathcal{K}}_{+}&:=\hat{\mathcal{K}}_{2}+\frac{1}{2}\hat{\mathcal{K}}_{3}-\biggl{(}\frac{1}{2}\gamma_{2}-2\gamma_{4}\biggr{)}\hat{\mathcal{K}}_{1}^{2},\\ \hat{\mathcal{K}}_{-}&:=\hat{\mathcal{K}}_{2}-\frac{1}{2}\hat{\mathcal{K}}_{3}-\biggl{(}\frac{1}{2}\gamma_{2}-2\gamma_{4}\biggr{)}\hat{\mathcal{K}}_{1}^{2},\end{split}$ (5.1)
and close on the commutation relations in (2.9) with
which further factorizes as
$\begin{split}\Phi(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})&=\Phi_{1}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})\Phi_{2}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}}),\qquad\bigl{[}\Phi_{1}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}}),\Phi_{2}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})\bigr{]}=0,\\ \Phi_{1}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})&=\frac{1}{4}\bigl{(}\hat{\mathcal{H}}_{4}-2{\rm i}\gamma_{1}\hat{\mathcal{K}}+4\gamma_{2}\hat{\mathcal{K}}^{2}+8{\rm i}\gamma_{3}\hat{\mathcal{K}}^{3}-16\gamma_{4}\hat{\mathcal{K}}^{4}\bigr{)},\\[2.0pt] \Phi_{2}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})&=\hat{\mathcal{H}}_{4}-2({\rm i}\gamma_{1}-2\gamma_{2}-4{\rm i}\gamma_{3}+8\gamma_{4}){\rm Id}+2({\rm i}\gamma_{1}-4\gamma_{2}-12{\rm i}\gamma_{3}+32\gamma_{4})\hat{\mathcal{K}}\\[2.0pt] &\qquad+4(\gamma_{2}+6{\rm i}\gamma_{3}-24\gamma_{4})\hat{\mathcal{K}}^{2}-8({\rm i}\gamma_{3}-8\gamma_{4})\hat{\mathcal{K}}^{3}-16\gamma_{4}\hat{\mathcal{K}}^{4}\\[2.0pt] &=\hat{\mathcal{H}}_{4}+2{\rm i}\gamma_{1}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}+4\gamma_{2}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{2}-8{\rm i}\gamma_{3}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{3}-16\gamma_{4}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{4}.\end{split}$ (5.2)
5.1 Spectrum of the quartic system
We introduce the finite-dimensional representation of (2.12) into the expressions (5.2), obtaining the structure function $\Phi=\Phi_{1}\Phi_{2}$ (2.13) as
$\begin{split}\Phi_{1}({B},E,u)&=\frac{1}{4}\bigl{(}E-2{\rm i}\gamma_{1}({B}+u)+4\gamma_{2}({B}+u)^{2}+8{\rm i}\gamma_{3}({B}+u)^{3}-16\gamma_{4}({B}+u)^{4}\bigr{)},\\[2.0pt] \Phi_{2}({B},E,u)&=E+2{\rm i}\gamma_{1}({B}+u-1)+4\gamma_{2}({B}+u-1)^{2}-8{\rm i}\gamma_{3}({B}+u-1)^{3}-16\gamma_{4}({B}+u-1)^{4}.\end{split}$ (5.3)
The spectrum $E$ of ^$\hat{\mathcal{H}}_{4}$ (2.4) is derived by imposing the two conditions (2.14) and is established as follows.
Proposition 3
(i) The set of equations (2.14) obtained from (5.3) leads to twelve types of solutions for the representation dependent constant $u=u({n})$ and the spectrum $E=E({n})$ of the Hamiltonian ^$\hat{\mathcal{H}}_{4}$ (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$. They provide the structure function $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}=\{1,\dots,{n}\}$ through (5.3).
(ii) Among all these solutions, there are only two types such that the limits $\gamma_{4}\to 0$ and $\gamma_{3}\to 0$ are simultaneously well-defined for any value of $\gamma_{1}$ and $\gamma_{2}$, namely
@@ 5.1 Spectrum of the quartic system
The ten excluded solutions exhibit divergencies, similarly to the solution (4.2) for the cubic Hamiltonian ^$\hat{\mathcal{H}}_{3}$, although the resulting expressions are obviously more complicated in this case (they involve both parameters $\gamma_{3}$ and $\gamma_{4}$ within the denominators). In this sense, the discarded solutions in propositions 2 and 3 can be regarded as ‘spurious’.
It is worth remarking that the results given in proposition 3 are well-defined for any value of the four parameters $\gamma_{k}$ $(k=1,\dots,4)$. As expected, they lead exactly to the results presented in proposition 2 for $\gamma_{4}=0$ and, then, to those characterized by proposition 1 when $\gamma_{3}=0$. In addition, it is possible to set $\gamma_{3}=0$ in the solutions (5.4), keeping the parameter $\gamma_{4}$, which determines a superintegrable quartic perturbation of the initial Hamiltonian ^$\hat{\mathcal{H}}_{2}$.
We observe that the expressions (5.4) take real values whenever the odd coefficients $\gamma_{1}$ and $\gamma_{3}$ are pure imaginary numbers, while the even ones $\gamma_{2}$ and $\gamma_{4}$ take real values. This difference between odd/even parameters has already been studied in the classical picture [3]. Therefore, by taking into account (4.3), let us denote
obtaining two types of spectra for ^$\hat{\mathcal{H}}_{4}$ given by
to be compared with (4.4). In addition, if we set $\alpha=-1$ and $\beta=-2$ and interpret ${n}$ as the principal quantum number $\mathfrak{n}$, we find two possible superintegrable perturbations of the spectrum (3.8) of the Zernike Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ in (1.1) depending on two real parameters $\mu$ and $\nu$, and given by
6 Higher-order quantum Zernike Hamiltonians: Conjectures
Although the quantum Hamiltonian ^$\hat{\mathcal{H}}_{N}$ has been analyzed in detail for the values $N<5$, a closed explicit expression for the higher-order quantum symmetries ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$ for arbitrary values of $N$ has not been found, due to the exponentially increasing computational difficulties. However, based on the results for the values analyzed, the form of the structure function $\Phi(\hat{\mathcal{H}}_{N},\hat{\mathcal{K}})$ in (2.9) along with its factorizing terms (2.10) can be inferred, allowing us to deduce the two types I and II of solutions for the spectrum that generalize those in (5.4) and preserve both the superintegrability and the properties concerning the limits $\gamma_{k}\rightarrow 0$ for $k\geq 3$. This extrapolation is presented in form of conjectures, and to illustrate their validity, we summarize the case for $N=5$, which has been explicitly solved, and indicate some of the computational obstructions to be expected for large values of $N$.
Conjecture 1
(i) For any $N$ and any value of the coefficients $\gamma_{k}$ ($1\leq k\leq N)$, the quantum Hamiltonian ^$\hat{\mathcal{H}}_{N}$ (2.3) admits three quantum symmetries corresponding to the quantum angular momentum operator $\hat{\mathcal{C}}$ in (2.5) and two operators ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$ (2.6) which are quadratic in the momenta for $N=1,2$ and of $N^{{\rm th}}$-order for $N>2$.
(ii) The sets {^$\{\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}_{N}\}$ and {^$\{\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}^{\prime}_{N}\}$ are formed by three algebraically independent operators, hence ^$\hat{\mathcal{H}}_{N}$ determines a superintegrable system.
(iii) The set $\{\hat{\mathcal{C}},{\hat{I}}_{N},{\hat{I}}^{\prime}_{N}\}$ leads to the quantum symmetries $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ in (2.8) and, in consequence, it is possible to define the number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ verifying the commutation relations (2.9) with the structure function $\Phi=\Phi_{1}\Phi_{2}$ (2.10), whose factorizing terms are given by
Observe that $[\Phi_{1},\Phi_{2}]=0$ and that the operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ should span a $(2N-1)^{{\rm th}}$-order polynomial symmetry algebra of Higgs-type $\mathfrak{sl}^{(2N-1)}(2,\mathbb{R})$ determined by the commutator $[\hat{\mathcal{K}}_{-},\hat{\mathcal{K}}_{+}]$.
@@ 6 Higher-order quantum Zernike Hamiltonians: Conjectures
(ii) The set of equations (2.14) coming from (6.2) gives rise to two solutions for the representation dependent constant $u=u({n})$ and the spectrum $E=E({n})$ of the Hamiltonian ^$\hat{\mathcal{H}}_{N}$ (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$, that hold for any values of the coefficients $\gamma_{k}$. From these solutions the structure function $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})=\Phi_{1}({B},{n})\Phi_{2}({B},{n})$ for ${B}=\{1,\dots,{n}\}$ is obtained using (6.2). The resulting expressions read as

### v2

@@ 4 The cubic quantum Zernike Hamiltonian
In this section we study the quantum symmetries and deduce the possible spectra for the cubic Hamiltonian $\hat{\mathcal{H}}_{3}$ (2.3) depending on three arbitrary parameters $\gamma_{1}$, $\gamma_{2}$ and $\gamma_{3}$. It is worthy to be observed that the cubic $\gamma_{3}$-term in $\hat{\mathcal{H}}_{3}$ can be interpreted as a superintegrable perturbation of the quantum Zernike Hamiltonian $\hat{\mathcal{H}}_{\rm Zk}$ in (2.2). Indeed, we require that all the expressions for $\hat{\mathcal{H}}_{2}$ in (2.3) obtained in the previous section are recovered whenever $\gamma_{3}$ is put equal to zero.
The two quantum symmetries (2.6) of $\hat{\mathcal{H}}_{3}$, that are of third-order in the momenta, are obtained after some computation, and adopt the form
@@
[…]
@@
The symmetry algebra of the cubic quantum Hamiltonian $\hat{\mathcal{H}}_{3}$ hence corresponds to a fifth-order polynomial Higgs-type algebra $\mathfrak{sl}^{(5)}(2,\mathbb{R})$ (see [4] for the classical picture).
@@ 4.1 Spectrum of the cubic system
(i) The set of equations (2.14) coming from (4.1) yields the corresponding solutions for the representation-dependent constant u=u(n)u=u({n}) and the spectrum E=E(n)E=E({n}) of the Hamiltonian $\hat{\mathcal{H}}_{3}$ in (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$. They lead to the structure function $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}=\{1,\dots,{n}\}$ via the factorization (4.1).
(ii) The requirement that the solutions are well-defined for any values of $\gamma_{1}$ and $\gamma_{2}$ for $\gamma_{3}=0$ reduces the set of solutions to only two, given by
@@
To illustrate the situation with the discarded solutions for the limit $\gamma_{3}\to 0$, let us consider, for instance, one of them given by
@@
The two consistent types of solutions given in proposition 2 constitute a cubic extension of those presented in proposition 1 and displayed in Table 1. We observe that types III and IV do not admit such a cubic generalization, implying that they are specific solutions for the quadratic Hamiltonian (2.3) with $N=2$ (although they are not well-defined if $\gamma_{2}=0$). The spectrum corresponding to types I and II takes real values whenever the coefficients $\gamma_{1}$ and $\gamma_{3}$ are pure imaginary numbers, while $\gamma_{2}$ is real. By considering (1.9), as before, we denote
@@
to be compared with (3.7). Furthermore, if we set $\alpha=-1$ and $\beta=-2$ and identify n{n} with the principal quantum number $\mathfrak{n}$, we obtain two possible superintegrable perturbations of the spectrum (3.8) of the Zernike Hamiltonian $\hat{\mathcal{H}}_{\rm Zk}$ in (1.1), reading as
@@ 4.2 Perturbations of the Euclidean, spherical and hyperbolic oscillators
Following the analysis in section 3.2, we can interpret the results of proposition 2 as superintegrable cubic perturbations of the isotropic oscillator on $\mathbf{E}^{2}$, $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$. To this end, let us consider the solution of type I (4.4) (eliminating the index ‘I’), again setting $\beta=-2$ and $\alpha=-\kappa$, but now keeping an arbitrary real parameter $\mu$, which characterizes the perturbation. Thus, we have six possibilities, according to the curvature $\kappa$ and the two signs of $\mu$.
For the flat Euclidean oscillator with $\kappa=0$, we start with a negative value for the perturbation parameter $\mu=-|\mu|$, obtaining that
@@
Hence, the spectrum is cubic in $n$, unbounded and the energy between two consecutive states grows quadratically with n{n}. It is worth noting that the behaviour of the spectrum (4.5) is quite similar to that of the spherical or Higgs oscillator (3.11) and (3.12), so we can say that it is a cubic spherical perturbation of the Euclidean oscillator. In figure 3 we plot the first energies (4.5) for three negative values of $\mu$ together with those of the non-perturbed Euclidean oscillator (3.10) (to be compared with figure 1).
@@
We require a positive spectrum and positive differences of two consecutive energies. The condition $E(n)>0$ implies that $\mu n^{2}<2$, giving rise to a limit value for the perturbation parameter: $\mu<2$. The second requirement yields a maximum value $n_{\rm max}$ for the positive integer $n\geq 1$, similarly to the hyperbolic oscillator (see (3.15)):
$E(n_{\rm max})-E(n_{\rm max}-1)$ $=2-\mu\big(3(n_{\rm max}-1)^{2}+3(n_{\rm max}-1)+1\big)$
$=2-\mu\big(3n_{\rm max}^{2}-3n_{\rm max}+1\big)>0.$
From this, we find that $n_{\rm max}$ is the greatest integer number that satisfies the following constraint
$1\leq n_{\rm max}<\frac{1}{2}+\frac{1}{6}\sqrt{\frac{24}{\mu}-3}\,.$ (4.7)
Note that if we consider the ground state $n_{\rm max}=1$, we recover that $\mu<2$. In this respect, we can say that (4.6) and (4.7) determine a cubic hyperbolic perturbation of the Euclidean oscillator that leads to a positive finite discrete spectrum with ${n}\in\{1,\dots,n_{\rm max}\}$. We represent the discrete spectrum for three values of $n_{\rm max}$ (so for $\mu$) in figure 4, which are lower than the Euclidean ones (3.10), and worthy of comparison with figure 2.
@@
Figure 4: The finite discrete spectrum (4.6) of the cubic hyperbolic perturbation of the oscillator on $\mathbf{E}^{2}$ for three values $\mu\in\{0.06,\,0.03,\,0.015\}$ corresponding to $n_{\rm max}\in\{3,\,5,\,7\}$ (4.7) starting from the lower dots, respectively, together with the Euclidean spectrum (3.10) with $\mu=0$ in the highest values.
Now we consider the sphere $\mathbf{S}^{2}$ with positive curvature $\kappa>0$ and a negative parameter $\mu=-|\mu|$. We arrive at the expression
@@
$E(n)=2n+\kappa n^{2}-\mu n^{3},\qquad E(n+1)-E(n)=2+\kappa(2n+1)-\mu(3n^{2}+3n+1).$ (4.8)
The requirement to obtain a positive spectrum leads to the restriction $\mu<2+\kappa$, while imposing positive differences of two consecutive energies implies that there is a maximum value $n_{\rm max}$ for the positive integer $n\geq 1$ according to the inequality
$E(n_{\rm max})-E(n_{\rm max}-1)=2+\kappa(2n_{\rm max}-1)-\mu\big(3n_{\rm max}^{2}-3n_{\rm max}+1\big)>0,$
thus arriving at
$1\leq n_{\rm max}<\frac{1}{2}+\frac{\kappa}{3\mu}+\frac{1}{6}\sqrt{\frac{4\kappa^{2}+24\mu}{\mu^{2}}-3}\ .$ (4.9)
Notice that if $n_{\rm max}=1$, then we recover that $\mu<2+\kappa$.
@@
Likewise, both types of perturbations can be applied to the oscillator on the hyperbolic space $\mathbf{H}^{2}$ with negative curvature $\kappa=-|\kappa|<0$ (see (3.13)–(3.15)); these are the spherical perturbation for $\mu=-|\mu|<0$ and the hyperbolic perturbation for $\mu>0$. A similar analysis to the one previously performed for the sphere $\mathbf{S}^{2}$ would be necessary for each type of perturbation. However, the results are quite cumbersome, as there are several possibilities for $n_{\rm max}$ depending on $|\kappa|$ and $\mu$. For brevity in the exposition, these explicit computations are omitted.
Figure 5: The finite discrete spectrum (4.8) of the cubic hyperbolic perturbation of the spherical oscillator on $\mathbf{S}^{2}$ with $\kappa=+1$ for $\mu\in\{0.2,\,0.12,\,0.1,\,0.07\}$ corresponding to $n_{\rm max}\in\{4,\,6,\,8,\,10\}$ (4.9) starting from the lower dots.
5 Higher-order quantum Zernike Hamiltonians: Conjectures
@@
Conjecture 1
(i) For any $N\geq 1$ and any value of the coefficients $\gamma_{k}$ ($1\leq k\leq N)$, the quantum Hamiltonian $\hat{\mathcal{H}}_{N}$ (2.3) admits three quantum symmetries corresponding to the quantum angular momentum operator $\hat{\mathcal{C}}$ in (2.5) and two operators ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$ (2.6), which are quadratic in the momenta for $N=1,2$ and of $N^{{\rm th}}$-order for $N>2$.
(ii) The sets $\big(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}_{N}\big)$ and $\big(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}^{\prime}_{N}\big)$ are formed by three algebraically independent operators, hence $\hat{\mathcal{H}}_{N}$ determines a superintegrable system.
@@
(ii) The set of equations (2.14) coming from (5.2) gives rise to two solutions for the representation dependent constant u=u(n)u=u({n}) and the spectrum E=E(n)E=E({n}) of the Hamiltonian $\hat{\mathcal{H}}_{N}$ (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$, that hold for any values of the coefficients $\gamma_{k}$. From these solutions the structure function $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})=\Phi_{1}({B},{n})\Phi_{2}({B},{n})$ for ${B}=\{1,\dots,{n}\}$ is obtained using (5.2). The resulting expressions read as